# Tubes in sub-Riemannian geometry and a Weyl’s invariance result for curves in the Heisenberg groups

### v2

@@ 5. Polynomial character of iterated divergences
To deal with the last item in (110), consider the bracket relations $[Y_{\alpha},Y_{\beta}]=c^{\varrho}_{\alpha\beta}Y_{\varrho}$, where $c^{\varrho}_{\alpha\beta}:\mathcal{O}\to\mathbb{R}$ are smooth (or real-analytic) We obtain the following distributional identity:
(112) $\begin{split}\nabla\delta(Y_{i}Y_{j}\delta)&=a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)(Y_{\beta}Y_{i}Y_{j}\delta)\\ &=a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)\left[(Y_{i}Y_{\beta}Y_{j}\delta)+c^{\varrho}_{\beta i}(Y_{\varrho}Y_{j}\delta)\right]\\ &=a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)(Y_{i}Y_{j}Y_{\beta}\delta)+a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)\left[Y_{i}\left(c^{\varrho}_{\beta j}(Y_{\varrho}\delta)\right)+c^{\varrho}_{\beta i}(Y_{\varrho}Y_{j}\delta)\right].\end{split}$
@@
The coefficient $m-1$ in the last variable of (116) is a useful notation to recall that, if $m=1$, there is no dependence on that variable.
@@
It is well-known that the group of isometries (fixing the origin) of $\mathbb{H}$ consists in all group automorphisms that act by orthogonal transformations on the distribution, [Ham90, LDO16]. We refer to [LR17, Sec. 6] for the explicit description we employ here. It is isomorphic to $\mathrm{SO}(2)$, and any isometry is determined by its action on the distribution at the origin. Let $\phi_{M}:\mathbb{H}\to\mathbb{H}$ be such an isometry for some $M\in\mathrm{SO}(2)$. The corresponding action on a left-invariant oriented orthonormal frame $X_{1},X_{2}$ is
@@ 6. Weyl’s invariance for curves in the Heisenberg groups
(131) $[X_{i},X_{j}]=J_{ij}X_{0},\qquad J:=\begin{pmatrix}0&\mathbbm{1}\\ -\mathbbm{1}&0\end{pmatrix},\qquad\forall\,i,j=1,\dots,2d,$
where $X_{0}=\partial_{z}$ is the Reeb vector field. Note that $[X_{0},X_{j}]=0$ for all $j=1,\dots,2d$.
@@
Fix $t\in[0,L]$. Up to applying a left-translation (which is an isometry), we may assume that $\gamma_{1}(t)=\gamma_{2}(t)=e$ is the Heisenberg group identity. Let $\Phi$ be a sub-Riemannian isometry such that $\Phi_{*}$ sends $W_{1}|_{e}$ to $W_{2}|_{e}$ and $X_{0}$ to ±$\pm X_{0}$. (Such an isometry exists by the structure of Heisenberg isometries, see e.g. [LR17, Sec. 6.1].) Then, by construction
@@
For $V\in T_{q}\Gamma$, we define a $(n-1)$-form on the $(n-1)$-dimensional submanifold $A_{q}\Gamma\subset A\Gamma$ (the fiber at $q$) by the position
@@
The integration of the $(n-1)$-form (142) on the disk of radius $r$ in $A_{q}\Gamma$ yields a one-parameter family of $1$-forms on $\Gamma$, denoted by $\mu_{\Gamma}^{r}$, defined as follows:
@@
(Note that $I=i^{*}$, where $i:M\to M$ is $i(x,y,z)=(-x,-y,-z)$. If $i$ were a sub-Riemannian isometry, then we would have $i\circ E=E\circ i^{-1*}=E\circ I$, and (163) would follow immediately, without the need of restriction to $T_{q}^{*}M$. However this is not the case and we have to argue differently.) By left-invariance of $E$ and $\mu$ it is sufficient to prove (163) at $q=e$ (the identity of the Heisenberg group). We use coordinates $(x,z)\in\mathbb{R}^{2d}\times\mathbb{R}$ on $M$, and dual coordinates $(p_{x},p_{z})\in\mathbb{R}^{2d}\times\mathbb{R}$ on each fiber $T^{*}_{e}M$, in terms of which it holds
@@
where, as computed e.g. in [Riz16, Lemma 15], $J(p_{x},p_{z})$ is the Jacobian determinant

### v3

@@ 5. Polynomial character of iterated divergences
To deal with the last item in (111), consider the bracket relations $Y_{\alpha},Y_{\beta}]=c^{\varrho}_{\alpha\beta}Y_{\varrho}$, where $c^{\varrho}_{\alpha\beta}:\mathcal{O}\to\mathbb{R}$ are smooth (or real-analytic) We obtain the following distributional identity:
$\begin{split}\nabla\delta(Y_{i}Y_{j}\delta)&=a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)(Y_{\beta}Y_{i}Y_{j}\delta)\\ &=a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)\left[(Y_{i}Y_{\beta}Y_{j}\delta)+c^{\varrho}_{\beta i}(Y_{\varrho}Y_{j}\delta)\right]\\ &=a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)(Y_{i}Y_{j}Y_{\beta}\delta)+a^{\alpha}_{h}a^{\beta}_{h}(Y_{\alpha}\delta)\left[Y_{i}\left(c^{\varrho}_{\beta j}(Y_{\varrho}\delta)\right)+c^{\varrho}_{\beta i}(Y_{\varrho}Y_{j}\delta)\right].\end{split}$
@@
The coefficient m1m-1 in the last variable of (116) is a useful notation to recall that, if $m=1$, there is no dependence on that variable.
@@
It is well-known that the group of isometries (fixing the origin) of $\mathbb{H}$ consists in all group automorphisms that act by orthogonal transformations on the distribution, [15, 18]. We refer to [20, Sec. 6] for the explicit description we employ here. It is isomorphic to $\mathrm{SO}(2)$, and any isometry is determined by its action on the distribution at the origin. Let $\phi_{M}:\mathbb{H}\to\mathbb{H}$ be such an isometry for some $M\in\mathrm{SO}(2)$. The corresponding action on a left-invariant oriented orthonormal frame $X_{1},X_{2}$ is
@@ 6. Weyl’s invariance for curves in the Heisenberg groups
(131) $X_{i},X_{j}]=J_{ij}X_{0},\qquad J:=\begin{pmatrix}0&\mathbbm{1}\\ -\mathbbm{1}&0\end{pmatrix},\qquad\forall\,i,j=1,\dots,2d,$
where $X_{0}=\partial_{z}$ is the Reeb vector field. Note that $X_{0},X_{j}]=0$ for all $j=1,\dots,2d$.
@@
Fix $t\in[0,L]$. Up to applying a left-translation (which is an isometry), we may assume that $\gamma_{1}(t)=\gamma_{2}(t)=e$ is the Heisenberg group identity. Let $\Phi$ be a sub-Riemannian isometry such that $\Phi_{*}$ sends $W_{1}|_{e}$ to $W_{2}|_{e}$ and $X_{0}$ to $\pm X_{0}$. (Such an isometry exists by the structure of Heisenberg isometries, see e.g. [20, Sec. 6.1].) Then, by construction
@@
For $V\in T_{q}\Gamma$, we define a (n1)(n-1)-form on the (n1)(n-1)-dimensional submanifold $A_{q}\Gamma\subset A\Gamma$ (the fiber at $q$) by the position
@@
The integration of the (n1)(n-1)-form (142) on the disk of radius $r$ in $A_{q}\Gamma$ yields a one-parameter family of $1$-forms on $\Gamma$, denoted by $\mu_{\Gamma}^{r}$, defined as follows:
@@
(Note that $I=i^{*}$, where $i:M\to M$ is i(x,y,z)=(x,y,z)i(x,y,z)=(-x,-y,-z). If $i$ were a sub-Riemannian isometry, then we would have $i\circ E=E\circ i^{-1*}=E\circ I$, and (163) would follow immediately, without the need of restriction to $T_{q}^{*}M$. However this is not the case and we have to argue differently.) By left-invariance of $E$ and $\mu$ it is sufficient to prove (163) at $q=e$ (the identity of the Heisenberg group). We use coordinates $(x,z)\in\mathbb{R}^{2d}\times\mathbb{R}$ on $M$, and dual coordinates $(p_{x},p_{z})\in\mathbb{R}^{2d}\times\mathbb{R}$ on each fiber $T^{*}_{e}M$, in terms of which it holds
@@
where, as computed e.g. in [23, Lemma 15], $J(p_{x},p_{z})$ is the Jacobian determinant